# No time to train! Training-Free Reference-Based Instance Segmentation

### v2

@@ 1 Introduction
Figure 1: Cross-domain 1-shot segmentation results using our training-free method on CD-FSOD benchmark. Our method directly evaluates on diverse datasets without any fine-tuning, using frozen SAMv2 and DINOv2 models. The reference set contains a single example image per class. The model then segments the entire target dataset based on the reference set. Results show: (1) generalization capabilities to out-of-distribution domains (e.g., underwater images, cartoons, microscopic textures); (2) state-of-the-art performance in 1-shot segmentation without training or domain adaptation; (3) limitations in cases with ambiguous annotations or highly similar classes (e.g., “harbor” vs. “ships” in DIOR). Best viewed when zoomed in. Ablation studies further investigate the variance associated with the selection of reference images. Refer to the Supplementary Material for more visualisations per dataset.
@@
It is well understood that collecting large-scale annotations for segmentation tasks is a costly and time-consuming process [2, 57, 17]. Recent advances in promptable segmentation frameworks [89, 73, 29, 74, 56, 44], epitomised by the Segment Anything Model (SAM) [33, 61], have significantly reduced manual effort by enabling high-quality mask generation using simple geometric prompts such as points, boxes or rough sketches. While this represents a substantial advancement in reducing manual effort, these masks lack semantic awareness [64, 9, 21, 27] and require either manual intervention or complex, domain-specific prompt-generation pipelines to function autonomously (e.g., medical imaging [48, 37, 84, 83], agriculture [5, 69], remote sensing [55, 49, 70]). Relying on manual prompts for each image limits their scalability (especially for large datasets or scenarios requiring automatic processing), while relying on domain-constrained automated pipelines restricts the ability to generalise to cross-domain scenarios.
@@
Reference-based instance segmentation [53, 15, 19] offers a promising solution to this challenge by using a small set of annotated reference images to guide the segmentation of a large set of target images. This idea has the potential to enable cheap, quick and automatic annotation of datasets where such labelling is expensive, time-consuming and requires expertise knowledge [2, 57]. Unlike slow manual prompting [24], using reference images can incorporate semantic understanding directly from examples, and is thus well-suited for automated segmentation tasks. Despite promising results, we observe that existing reference-based segmentation methods often require fine-tuning on novel classes and this raises a set of well understood concerns that include task-specific data requirements, overfitting and domain shift. We conjecture that a prospective alternative approach to guiding reference-based instance segmentation involves reusing the general purpose capabilities of vision foundation models [4, 60, 54, 33, 61].
@@
Several works [46, 67, 74] have attempted to combine pretrained models for reference-based segmentation, e.g. Matcher [46] which integrates DINOv2 with SAM for semantic segmentation tasks. However, these methods face several limitations. Firstly, they rely on computationally expensive distance metrics (Earth Mover’s Distance), and complex thresholding mechanisms, which significantly slow down inference. Secondly, they are not suited for instance-level segmentation tasks, struggling with fine-grained discrimination in complex multi-object scenes. In fact, the instance segmentation setting presents unique challenges—how do we handle occlusions, scale variations, ambiguous object boundaries and varying image quality, all with just a few reference images—and they should be tackled thoughtfully. Effectively combining foundation models, without significant finetuning, remains a large challenge [9], particularly when attempting to leverage generalisation capabilities of semantic ViT backbones (e.g. DINOv2), to achieve precise localisation [82].
We propose a training-free three-stage method: (1) constructing a memory bank of category-specific features, (2) refining feature representations via two-step aggregation, and (3) performing inference through feature matching and a novel semantic-aware soft merging strategy.
@@
This results in a training-free, high-performing framework that achieves significant performance gains on established datasets. Furthermore, our approach maintains its effectiveness across diverse domains with fixed hyperparameters, making it accessible for a wide range of applications. Our contributions can be summarised as:
@@
  • We introduce a novel three-stage framework for instance segmentation with vision foundation models, addressing key integration challenges through (1) memory bank construction, (2) two-step feature aggregation, and (3) feature matching with semantic-aware soft merging.
@@ 2 Related Work
Reference-based Instance Segmentation aims to segment individual objects within an image, distinguishing between instances of the same category [6, 39]. Traditional approaches, such as Mask R-CNN [22], use region proposals with convolutional networks to predict instance masks, while transformer-based models like DETR [3] and Mask2Former [7] integrate global context through self-attention mechanisms. These methods have demonstrated success in standard instance segmentation tasks by leveraging large, labeled datasets [42, 17]. Reference-based instance segmentation extends these tasks to handle novel categories with limited labeled examples. Early works [15, 53] adapted Mask R-CNN by introducing instance-level discriminative features [15] or uncertainty-guided bounding box prediction [53]. More recent works have unified segmentation tasks into in-context learning frameworks [73, 74, 29], involving an expensive pretraining phase on a wide range of segmentation tasks, including instance segmentation [73, 74] and using contrastive pretraining to integrate visual and textual prompts [29]. Despite these advancements, reference-based instance segmentation remains challenging due to the lack of labeled data, the complexity of multi-instance scenarios, limited generalisation across domains for specialist models, reference image ambiguity, and reliance on predefined class labels [79]. Additionally, reusing frozen backbones not originally pretrained for instance segmentation remains a challenge [82, 9]. Our method effectively reuses two existing frozen vision foundation models—none of which was trained for reference-based instance segmentation task— to tackle reference-based instance segmentation without additional training, while also generalising well to unusual domains.
Vision Foundation models have revolutionised computer vision by learning strong, pretrained representations, transferable across diverse tasks. CLIP [60] and DINO models [4, 54] exemplify this trend, using contrastive learning to align visual and textual representations, and learning robust image embeddings from unlabeled data, respectively. These models have been widely adopted for downstream tasks, including open-vocabulary detection [64, 21] and semantic segmentation [62, 88]. However, CLIP struggles with detailed spatial reasoning, while DINOv2, despite capturing fine-grained semantics, produces low-resolution feature maps. The Segment Anything Models (SAM, SAM2) [33, 61] are a notable addition to this category, trained on an extensive, category-agnostic dataset (SA-1B) [33]. While SAM excels in generating segmentation masks with minimal input (e.g., points, bounding boxes), it lacks inherent semantic understanding [64, 9, 21, 27]. Efforts to bridge this gap include pairing SAM with language models [34], diffusion models [87], or fine-tuning it on labeled datasets like COCO [42] and ADE20K [38]. However, these adaptations often result in complex pipelines or limited scalability. Furthermore, SAM’s semantic-agnostic nature poses limitations in scenarios requiring class differentiation at instance-level. Our work combines the complementary strengths of DINOv2 and SAM without requiring finetuning. By aggregating and matching features from multiple references, we enable high-precision, training-free instance segmentation, achieving state-of-the-art results on diverse few-shot benchmarks.
@@
Automatic Vision Prompting for SAM aims to construct automatic prompting pipelines to enhance SAM’s versatility in complex visual tasks, reducing its reliance on manual inputs. Training-free methods [81, 46] leverage feature-matching techniques but often rely on manually tuned thresholds, distance metrics and complex pipelines. Other approaches focus on learning prompts directly, such as spatial or semantic optimisation [28, 67, 26], but these methods still face challenges in multi-instance or semantically dense settings. Zero-shot methods [78, 65] introduce visual markers to guide attention during segmentation, but lack fine-grained precision. Recent efforts, including SEEM [89], have unified segmentation and recognition tasks through shared decoders, while SINE [45] tackles task ambiguity by disentangling segmentation tasks. Others have paired SAM with Stable Diffusion for open-vocabulary segmentation [87], enabling it to incorporate semantic cues. Similarly, LISA [34] utilises language-based instructions to adapt SAM for text-guided tasks. Despite these efforts, handling multi-instance scenarios and semantic ambiguity without training remains challenging. Our method differs by integrating SAM with DINOv2 without additional training or prompt optimisation. We introduce a multi-stage method (memory bank construction, representation aggregation, and semantic-aware feature matching) with fixed hyperparameters across all experimental settings, making it practical for a wide range of downstream applications and directly accessible to practitioners across diverse domains.
@@ 3.1 Preliminaries
Segment Anything Model (SAM) [33] is designed for promptable segmentation, that is, it responds to different types of geometric prompts to generate image segmentation masks. It consists of three main components: an image encoder, a prompt encoder, and a mask decoder. The image encoder is a pretrained Vision Transformer (ViT) [8] adapted for high-resolution inputs [40]. The prompt encoder handles both sparse (points, boxes, text) and dense (rough mask) prompts, which are encoded with positional encodings [68], learnt embeddings, and off-the-shelf text encoders. The mask decoder efficiently generates masks by using a modified Transformer decoder block [3] with self-attention and cross-attention on the prompts. For training, SAM uses a combination of focal loss [43] and dice loss [50].
DINOv2 Self-supervised pretrained vision encoder [54] is a self-supervised vision model designed to produce general-purpose visual features. It uses a discriminative self-supervised learning approach based on Vision Transformer (ViT) architectures [8]. DINOv2 makes use of teacher-student networks, incorporates Sinkhorn-Knopp normalisation, multi-crops strategy, separate projection heads for image-level and patch-level objectives, and additional regularisation techniques to stabilise and scale training [4, 85]. It is trained on a curated dataset (LVD-142M) of 142 million images using efficient training techniques (fast memory-efficient attention, stochastic depth, etc.). Its features are transferable across tasks and domains, making it a robust backbone for both global and local visual understanding tasks.
Reference-based instance segmentation aims to segment a target image by using a reference segmented image as an example (rather than using geometric prompts). More formally, given a reference image $I_{r}$ and its corresponding annotations $M^{i}_{r}$ (where $i$ denotes the different object categories), we use this data to segment the corresponding regions in the target image $I_{t}$ that also belong to category $i$.
@@ 3.2 Training-free method
The goal of our training-free method is to extract category-specific features from a set of annotated reference examples and use them to segment and classify instances in target images. Unlike methods that require model retraining, we employ a memory-based approach to store discriminative representations for object categories. It consists of three main stages: (1) constructing a memory bank from reference images, (2) refining these representations via two-stage feature aggregation, and (3) performing inference on the target images through feature matching and semantic-aware soft merging. Figure 2 illustrates the complete pipeline.
(1) Memory Bank Construction.
Given a set of reference images $\{I_{r}^{j}\}_{j=1}^{N_{r}}$ and their corresponding instance masks $\{M_{r}^{j,i}\}_{j=1}^{N_{r}}$ for category $i$, we extract dense feature maps $F_{r}^{j}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times d}$ using a pretrained frozen encoder $\mathcal{E}$ [^1] https://github.com/facebookresearch/dinov2, where $d$ is the feature dimension and $H^{\prime},W^{\prime}$ denote the spatial resolution of the feature map. The corresponding instance masks $M_{r}^{j,i}\in\{0,1\}^{H^{\prime}\times W^{\prime}}$ are resized to match this resolution.
For each category $i$, we store the masked features:
$\mathcal{F}_{r}^{j,i}=F_{r}^{j}\odot M_{r}^{j,i}$
where $\odot$ denotes element-wise multiplication. These category-wise feature sets are stored in a memory bank $\mathcal{M}_{i}$, which is synchronised across GPUs to ensure consistency in distributed settings.
(2) Two-stage feature aggregation.
To construct category prototypes, we first compute instance-wise feature representations, and then, aggregate them into class-wise prototypes.
@@
    Instance-wise prototypes:
    Each instance $k$ in reference image $I_{r}^{j}$ has its own prototype, computed by averaging the feature embeddings within its corresponding mask:
    $P_{r}^{j,k}=\frac{1}{\|M_{r}^{j,k}\|_{1}}\sum_{(u,v)}M_{r}^{j,k}(u,v)F_{r}^{j}(u,v)$
    where $P_{r}^{j,k}\in\mathbb{R}^{d}$ represents the mean feature representation of the $k$-th instance in image $I_{r}^{j}$.
@@
    Class-wise prototype:
    We compute the category prototype $P_{i}$ by averaging all instance-wise prototypes belonging to the same category $i$:
    $P_{i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{r}}\sum_{k\in\mathcal{K}_{i}^{j}}P_{r}^{j,k},$
    where $\mathcal{K}_{i}^{j}$ is the set of instances in image $I_{r}^{j}$ that belong to category $i$, and $N_{i}=\sum_{j=1}^{N_{r}}|\mathcal{K}_{i}^{j}|$ is the total number of instances belonging to category $i$. These class-wise prototypes $P_{i}$ are stored in the memory bank.
(3) Inference on Target Images.
For a target image $I_{t}$, we extract dense features $F_{t}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times d}$ using the same encoder $\mathcal{E}$. We use the frozen SAM model to generate $N_{m}$ candidate instance masks $\{M_{t}^{m}\}_{m=1}^{N_{m}}$, where $M_{t}^{m}\in\{0,1\}^{H^{\prime}\times W^{\prime}}$. Each mask $M_{t}^{m}$ is used to compute a feature representation via average pooling and L2 normalisation:
$P_{t}^{m}=\frac{1}{\|M_{t}^{m}\|_{1}}\sum_{(u,v)}M_{t}^{m}(u,v)F_{t}(u,v),\quad\hat{P}_{t}^{m}=\frac{P_{t}^{m}}{\|P_{t}^{m}\|_{2}}$
where $\hat{P}_{t}^{m}\in\mathbb{R}^{d}$ is the normalised mask feature.
To classify each candidate mask, we compute:
@@
    Feature Matching. We compute the cosine similarity between $\hat{P}_{t}^{m}$ and category prototypes $P_{i}$:
    $S_{t}^{m}=\max_{i}\left(\frac{\hat{P}_{t}^{m}\cdot P_{i}}{\|P_{i}\|_{2}}\right)$
    which provides the classification score $S_{t}^{m}$ for mask $M_{t}^{m}$.
@@
    Semantic-Aware Soft Merging. To handle overlapping predictions, we introduce a novel soft merging strategy. Given two masks $M_{t}^{m}$ and $M_{t}^{m^{\prime}}$ of the same category, we compute their intersection-over-self (IoS):
    $\text{IoS}(M_{t}^{m},M_{t}^{m^{\prime}})=\frac{\sum(M_{t}^{m}\cap M_{t}^{m^{\prime}})}{\sum M_{t}^{m}}$
    and weight it by feature similarity:
    $w_{m,m^{\prime}}=\frac{\hat{P}_{t}^{m}\cdot\hat{P}_{t}^{m^{\prime}}}{\|\hat{P}_{t}^{m^{\prime}}\|_{2}}$
    The final score for each mask is adjusted using a decay factor:
    $S_{t}^{m}\leftarrow S_{t}^{m}\cdot\sqrt{(1-\text{IoS}(M_{t}^{m},M_{t}^{m^{\prime}})w_{m,m^{\prime}}})$
    reducing redundant detections while preserving distinct instances that may partially overlap. Finally, we rank masks by their adjusted scores, and select the top-$K$ predictions as the final output.

### v3

@@ 1 Introduction
Figure 1: Cross-domain 1-shot segmentation results using our training-free method on CD-FSOD benchmark. Our method directly evaluates on diverse datasets without any fine-tuning, using frozen SAMv2 and DINOv2 models. The reference set contains a single example image per class. The model then segments the entire target dataset based on the reference set. Results show: (1) generalization capabilities to out-of-distribution domains (e.g., underwater images, cartoons, microscopic textures); (2) state-of-the-art performance in 1-shot segmentation without training or domain adaptation; (3) limitations in cases with ambiguous annotations or highly similar classes (e.g., “harbor” vs. “ships” in DIOR). Best viewed when zoomed in. Ablation studies further investigate the variance associated with the selection of reference images. See Appendix A.8 for more visualisations.
@@
It is well understood that collecting large-scale annotations for segmentation tasks is a costly and time-consuming process (Benenson et al., 2019; Papadopoulos et al., 2021; Gupta et al., 2019). Recent advances in promptable segmentation frameworks (Zou et al., 2023; Wang et al., 2023b; Jiang et al., 2024; Wang et al., 2023c; Pan et al., 2023; Liu et al., 2024a), epitomised by the Segment Anything Model (SAM) (Kirillov et al., 2023; Ravi et al., 2024), have significantly reduced manual effort by enabling high-quality mask generation using simple geometric prompts such as points, boxes or rough sketches. While this represents a substantial advancement in reducing manual effort, these masks lack semantic awareness (Shin et al., 2024; Espinosa et al., 2024; Han et al., 2023; Ji et al., 2024) and require either manual intervention or complex, domain-specific prompt-generation pipelines to function autonomously e.g. in medical imaging (Ma et al., 2024a; Lei et al., 2023; Zhang et al., 2024c; Zhang and Jiao, 2023), agriculture (Carraro et al., 2023; Tripathy et al., 2024), remote sensing (Osco et al., 2023; Ma et al., 2024b; Wang et al., 2023a)). Relying on manual prompts for each image limits their scalability (especially for large datasets or scenarios requiring automatic processing), while relying on domain-constrained automated pipelines restricts the ability to generalise to cross-domain scenarios.
@@
Reference-based instance segmentation (Nguyen and Todorovic, 2022; Ganea et al., 2021; Han et al., 2022a) offers a promising solution to this challenge by using a small set of annotated reference images to guide the segmentation of a large set of target images. This idea has the potential to enable cheap, quick and automatic annotation of datasets where such labelling is expensive, time-consuming and requires expertise knowledge (Benenson et al., 2019; Papadopoulos et al., 2021). Unlike slow manual prompting (Hu et al., 2024), using reference images can incorporate semantic understanding directly from examples, and is thus well-suited for automated segmentation tasks. Despite promising results, we observe that existing reference-based segmentation methods often require fine-tuning on novel classes and this raises a set of well understood concerns that include task-specific data requirements, overfitting and domain shift. We conjecture that a prospective alternative approach to guiding reference-based instance segmentation involves reusing the general purpose capabilities of vision foundation models (Caron et al., 2021; Radford et al., 2021; Oquab et al., 2024; Kirillov et al., 2023; Ravi et al., 2024).
@@
Several works (Liu et al., 2024c; Sun et al., 2024; Wang et al., 2023c) have attempted to combine pretrained models for reference-based segmentation, e.g. Matcher (Liu et al., 2024c) which integrates DINOv2 with SAM for semantic segmentation tasks. However, these methods face several limitations. Firstly, they rely on computationally expensive distance metrics (Earth Mover’s Distance), and complex thresholding mechanisms, which significantly slow down inference. Secondly, they are not suited for instance-level segmentation tasks, struggling with fine-grained discrimination in complex multi-object scenes. In fact, the instance segmentation setting presents unique challenges—how do we handle occlusions, scale variations, ambiguous object boundaries and varying image quality, all with just a few reference images—and they should be tackled thoughtfully. Effectively combining foundation models, without significant finetuning, remains a large challenge (Espinosa et al., 2024), particularly when attempting to leverage generalisation capabilities of semantic ViT backbones (e.g. DINOv2), to achieve precise localisation (Zhang et al., 2024b). These observations highlight that naively composing existing foundation models is insufficient for effective instance-level matching; achieving strong performance requires a dedicated system-level design, which motivates the training-free method we develop in this work. Code is made available.
@@
We propose a training-free three-stage method: (1) constructing a memory bank of category-specific features, (2) refining feature representations via two-step aggregation, and (3) performing inference through feature matching and a novel semantic-aware soft merging strategy. This results in a training-free, high-performing framework that achieves significant gains on established datasets. Furthermore, our approach maintains its effectiveness across diverse domains with fixed hyperparameters, making it accessible for a wide range of applications. Our contributions are:
@@
  • We introduce a novel three-stage framework for instance segmentation with vision foundation models, addressing key integration challenges with (1) memory bank construction, (2) two-step feature aggregation, and (3) feature matching with semantic-aware soft merging.
@@ 2 Related Work
Reference-based Instance Segmentation aims to segment individual objects within an image, distinguishing between instances of the same category (Chen et al., 2019; Li et al., 2017). Traditional approaches, such as Mask R-CNN (He et al., 2017), use region proposals with convolutional networks to predict instance masks, while transformer-based models like DETR (Carion et al., 2020) and Mask2Former (Cheng et al., 2022) integrate global context through self-attention mechanisms. These methods have demonstrated success in standard instance segmentation tasks by leveraging large, labeled datasets (Lin et al., 2014; Gupta et al., 2019). Reference-based instance segmentation extends these tasks to handle novel categories with limited labeled examples. Early works (Ganea et al., 2021; Nguyen and Todorovic, 2022) adapted Mask R-CNN by introducing instance-level discriminative features (Ganea et al., 2021) or uncertainty-guided bounding box prediction (Nguyen and Todorovic, 2022). More recent works have unified segmentation tasks into in-context learning frameworks (Wang et al., 2023b, c; Jiang et al., 2024); (Ren et al., 2024), involving an expensive pretraining phase on a wide range of segmentation tasks, including instance segmentation (Wang et al., 2023b, c); (Ren et al., 2024) and using contrastive pretraining to integrate visual and textual prompts (Jiang et al., 2024). Despite these advancements, reference-based instance segmentation remains challenging due to the lack of labeled data, the complexity of multi-instance scenarios, limited generalisation across domains for specialist models, reference image ambiguity, and reliance on predefined class labels (Zang et al., 2021). Additionally, reusing frozen backbones not originally pretrained for instance segmentation remains a challenge (Zhang et al., 2024b; Espinosa et al., 2024). Our method effectively reuses two existing frozen vision foundation models—none of which was trained for reference-based instance segmentation task— to tackle reference-based instance segmentation without additional training, while also generalising to unusual domains.
Vision Foundation models have revolutionised computer vision by learning strong, pretrained representations, transferable across diverse tasks. CLIP (Radford et al., 2021) and DINO models (Caron et al., 2021; Oquab et al., 2024) exemplify this trend, using contrastive learning to align visual and textual representations, and learning robust image embeddings from unlabeled data, respectively. These models have been widely adopted for downstream tasks, including open-vocabulary detection (Shin et al., 2024; Han et al., 2023) and semantic segmentation (Salehi et al., 2023; Ziegler and Asano, 2022). However, CLIP struggles with detailed spatial reasoning, while DINOv2, despite capturing fine-grained semantics, produces low-resolution feature maps. The Segment Anything Models (SAM, SAM2) (Kirillov et al., 2023; Ravi et al., 2024) are a notable addition to this category, trained on an extensive, category-agnostic dataset (SA-1B) (Kirillov et al., 2023). While SAM excels in generating segmentation masks with minimal input (e.g., points, bounding boxes), it lacks inherent semantic understanding (Shin et al., 2024; Espinosa et al., 2024; Han et al., 2023; Ji et al., 2024). Efforts to bridge this gap include pairing SAM with language models (Lai et al., 2024), diffusion models (Zhu et al., 2024), or fine-tuning it on labeled datasets like COCO (Lin et al., 2014) and ADE20K (Li et al., 2024). However, these adaptations often result in complex pipelines or limited scalability. Furthermore, SAM’s semantic-agnostic nature poses limitations in scenarios requiring class differentiation at instance-level. Our work combines the complementary strengths of DINOv2 and SAM without requiring finetuning. By aggregating and matching features from multiple references, we enable high-precision, training-free instance segmentation, achieving state-of-the-art results on diverse few-shot benchmarks.
@@
Automatic Vision Prompting for SAM aims to construct automatic prompting pipelines to enhance SAM’s versatility in complex visual tasks, reducing its reliance on manual inputs. Training-free methods (Zhang et al., 2024a; Liu et al., 2024c) leverage feature-matching techniques but often rely on manually tuned thresholds, distance metrics and complex pipelines. Other approaches focus on learning prompts directly, such as spatial or semantic optimisation (Jia et al., 2022; Sun et al., 2024; Huang et al., 2024), but these methods still face challenges in multi-instance or semantically dense settings. Zero-shot methods (Yao et al., 2024; Shtedritski et al., 2023) introduce visual markers to guide attention during segmentation, but lack fine-grained precision. Recent efforts, including SEEM (Zou et al., 2023), have unified segmentation and recognition tasks through shared decoders, while SINE (Liu et al., 2024b) tackles task ambiguity by disentangling segmentation tasks. Others have paired SAM with Stable Diffusion for open-vocabulary segmentation (Zhu et al., 2024), enabling it to incorporate semantic cues. Similarly, LISA (Lai et al., 2024) utilises language-based instructions to adapt SAM for text-guided tasks. Despite these efforts, handling multi-instance scenarios and semantic ambiguity without training remains challenging. Our method differs by integrating SAM with DINOv2 without additional training or prompt optimisation. We introduce a multi-stage method (memory bank construction, representation aggregation, and semantic-aware feature matching) with fixed hyperparameters across all experimental settings, making it practical for a wide range of downstream applications and directly accessible to practitioners across domains.
@@ 3.1 Preliminaries
Segment Anything Model (SAM) (Kirillov et al., 2023) is designed for promptable segmentation, that is, it responds to different types of geometric prompts to generate segmentation masks. It consists of three components: an image encoder, a prompt encoder, and a mask decoder. The image encoder is a Vision Transformer (ViT) (Dosovitskiy et al., 2021) adapted for high-resolution inputs (Li et al., 2022a). The prompt encoder handles sparse (points, boxes, text) and dense (mask) prompts, which are encoded with positional encodings (Tancik et al., 2020), learnt embeddings, and off-the-shelf text encoders. The mask decoder generates masks by using a modified Transformer decoder block (Carion et al., 2020) with self-attention and cross-attention on the prompts. For training, SAM uses a combination of focal loss (Lin et al., 2017) and dice loss (Milletari et al., 2016).
DINOv2 vision encoder (Oquab et al., 2024) is a self-supervised ViT-based (Dosovitskiy et al., 2021) model designed to produce general-purpose visual features using a discriminative self-supervised learning approach. DINOv2 makes use of teacher-student networks, incorporates Sinkhorn-Knopp normalisation, multi-crops strategy, separate projection heads for image-level and patch-level objectives, and additional regularisation techniques to stabilise and scale training (Caron et al., 2021; Zhou et al., 2022a). It is trained on a curated dataset (LVD-142M) of 142 million images using efficient training techniques (memory-efficient attention, stochastic depth, etc.). Its features are transferable across tasks and domains, making it a robust backbone for global and local visual understanding tasks.
Reference-based instance segmentation aims to segment a target image by using a reference segmented image (rather than geometric prompts). Given a reference image $I_{r}$ and its annotations $M^{i}_{r}$ (where $i$ denotes different object categories), we use this data to segment the corresponding regions in the target image $I_{t}$ that also belong to category $i$.
@@ 3.2 Training-free method
The goal is to extract category-specific features from a set of annotated reference examples and use them to segment and classify instances in target images. Unlike methods that require model retraining, we employ a memory-based approach to store discriminative representations for object categories. It consists of three main stages (see Figure 2): (1) constructing a memory bank from reference images, (2) refining these representations via two-stage feature aggregation, and (3) performing inference on the target images through feature matching and semantic-aware soft merging.
(1) Memory Bank Construction. Given a set of reference images $\{I_{r}^{j}\}_{j=1}^{N_{r}}$ and their instance masks $\{M_{r}^{j,i}\}_{j=1}^{N_{r}}$ for category $i$, we extract dense feature maps $F_{r}^{j}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times d}$ using a frozen encoder $\mathcal{E}$ [^1] https://github.com/facebookresearch/dinov2, where $d$ is the feature dimension and $H^{\prime},W^{\prime}$ denote the spatial resolution of the feature map. Instance masks $M_{r}^{j,i}\in\{0,1\}^{H^{\prime}\times W^{\prime}}$ are resized to match this resolution. For each category $i$, we store the masked features $\mathcal{F}_{r}^{j,i}=F_{r}^{j}\odot M_{r}^{j,i}$ where $\odot$ denotes element-wise multiplication. These category-wise feature sets are stored in a memory bank $\mathcal{M}_{i}$, which is synchronised across GPUs.
(2) Two-stage feature aggregation. To construct category prototypes, we first compute instance-wise feature representations, and then, aggregate them into class-wise prototypes.
@@
    Instance-wise prototypes: Each instance $k$ in reference image $I_{r}^{j}$ has its own prototype, computed by averaging the feature embeddings within its corresponding mask: $P_{r}^{j,k}=\frac{1}{\|M_{r}^{j,k}\|_{1}}\sum_{(u,v)}M_{r}^{j,k}(u,v)F_{r}^{j}(u,v)$ where $P_{r}^{j,k}\in\mathbb{R}^{d}$ is the mean feature representation of the $k$-th instance in image $I_{r}^{j}$.
@@
    Class-wise prototype: We compute the category prototype $P_{i}$ by averaging all instance-wise prototypes belonging to the same category $i$: $P_{i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{r}}\sum_{k\in\mathcal{K}_{i}^{j}}P_{r}^{j,k},$ where $\mathcal{K}_{i}^{j}$ is the set of instances in image $I_{r}^{j}$ that belong to category $i$, and $N_{i}=\sum_{j=1}^{N_{r}}|\mathcal{K}_{i}^{j}|$ is the total number of instances belonging to category $i$. These class-wise prototypes $P_{i}$ are stored in the memory bank.
(3) Inference on Target Images. Given a target image $I_{t}$, we extract dense features $F_{t}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times d}$ using $\mathcal{E}$. We use SAM to generate $N_{m}$ candidate instance masks $\{M_{t}^{m}\}_{m=1}^{N_{m}}$, where $M_{t}^{m}\in\{0,1\}^{H^{\prime}\times W^{\prime}}$. Each mask $M_{t}^{m}$ is used to compute a feature representation $P_{t}^{m}$ via average pooling and L2-norm: $P_{t}^{m}=\frac{1}{\|M_{t}^{m}\|_{1}}\sum_{(u,v)}M_{t}^{m}(u,v)F_{t}(u,v),$ $\quad\hat{P}_{t}^{m}=\frac{P_{t}^{m}}{\|P_{t}^{m}\|_{2}}$ where $\hat{P}_{t}^{m}\in\mathbb{R}^{d}$.
To classify each candidate mask, we perform:
@@
    Feature Matching. We compute the cosine similarity between $\hat{P}_{t}^{m}$ and category prototypes $P_{i}$, which provides the classification score $S_{t}^{m}$ for mask $M_{t}^{m}$: $S_{t}^{m}=\max_{i}\left(\frac{\hat{P}_{t}^{m}\cdot P_{i}}{\|P_{i}\|_{2}}\right)$
@@
    Semantic-Aware Soft Merging. To handle overlapping predictions, we introduce a novel soft merging strategy. Given two masks $M_{t}^{m}$ and $M_{t}^{m^{\prime}}$ of the same category, we compute their intersection-over-self (IoS), and weight it by feature similarity.
    $\text{IoS}(M_{t}^{m},M_{t}^{m^{\prime}})\!=\!\frac{\sum(M_{t}^{m}\cap M_{t}^{m^{\prime}})}{\sum M_{t}^{m}}\quad w_{m,m^{\prime}}\!=\!\frac{\hat{P}_{t}^{m}\cdot\hat{P}_{t}^{m^{\prime}}}{\|\hat{P}_{t}^{m^{\prime}}\|_{2}}$
    The final score for each mask is adjusted using a decay factor $S_{t}^{m}\!\leftarrow\!S_{t}^{m}\!\cdot\!\sqrt{(1\!-\!\text{IoS}(M_{t}^{m},\,M_{t}^{m^{\prime}})w_{m,m^{\prime}}})$, reducing redundant detections while preserving distinct instances that may partially overlap. We rank masks by their adjusted scores, and select top-$K$ predictions.